OpenStudy (anonymous):

fan and medal A person leaves home at 8:00 a.m. and drives to a destination at a rate of 40 mph. The person returns at a rate of 25 mph and arrives at 2:30 p.m. How far was the destination? 65 miles 100 miles 130 miles

OpenStudy (radar):

The assumption is he immediately turned around at the destination???

OpenStudy (anonymous):

yes

OpenStudy (jjuden):

100 miles

OpenStudy (anonymous):

this is the only info they gave me

OpenStudy (radar):

surjithayer is providing the needed help.

OpenStudy (jjuden):

Total time 6.5 hr (Assuming no stopping time) D/r = t D/40mph + D/25mph = 6.5 65D = 6.5(40*25) D = 6.5(40*25)/65 D = 100mi

OpenStudy (anonymous):

thanks jjuden but im goin to see what the other person says to ....no offense to you

OpenStudy (anonymous):

not saying your wrong either

OpenStudy (anonymous):

Let x be the distance between home and destination. time taken to go from home to destination=x/40 hours time taken to go back from destination to home x/25 hours x/40+x/25=6.30hours=13/2 hours solve for x

OpenStudy (radar):

Not complaining about the 100 mi answer.

OpenStudy (anonymous):

\[x \frac{5+8 }{ 200 }=\frac{ 13 }{ 2 },x=\frac{ 13 }{ 2}*\frac{ 200 }{ 13 }=100m\]

OpenStudy (anonymous):

https://media.glynlyon.com/g_alg01_2012/10/m90659.gif could i use this graph

OpenStudy (j2lie):

I think it is 65 miles because you add 25 mph and 40 mph.

OpenStudy (radar):

Don't understand the table . time = 2. I believe it is 6.5 hr.

OpenStudy (j2lie):

Do you add or multiply?